Angles A and B are complementary angles. Determine the measures of angles A and B if angle A is four times the size of angle B.
Answer by nerdybill(7384)

Angles A and B are complementary angles. Determine the measures of angles A and B if angle A is four times the size of angle B.
.
Let x = measure of angle B
then
4x = measure of angle A
.
x + 4x = 90
5x = 90
x = 90/5
x = 18 degrees (angle B)
.
Angle A:
4x = 4(18) = 72 degrees